# Uncertainty Quantification on Graph Learning: A Survey

### v2

@@ IV Methods for Uncertainty Representation
TABLE II: Taxonomy for Uncertainty Quantification Methods
Purpose Uncertainty Representation (Sec. IV) Uncertainty Handling (Sec. V)
Methodology Bayesian✓ Direct Inference Bayesian✓ Representation Learning Calibration Conformal Prediction OOD Detection
Post-hoc
Uncertain Parameter
Distribution Free Input
Regression Task
Classification Task
Aleatoric and Epistemic
Uncertainty Separation
Accuracy Impact
Back Propagation
Computational Cost
High High Low Low Low
In the realm of UQ, Bayesian methods stand as a pivotal approach, offering a powerful framework for representing uncertainty. At the heart of Bayesian UQ is the application of Bayes’ Theorem, which provides a mathematical framework for updating our knowledge about a parameter $\boldsymbol{\theta}$ when we observe new data $D$. The theorem combines prior probability $P(\boldsymbol{\theta})$ with the likelihood $P(D|\boldsymbol{\theta})$ to produce the posterior probability $P(\boldsymbol{\theta}|D)$, normalized by the marginal likelihood $P(D)$:
@@
Advancing from the foundational Bayesian principles, we delve into three distinct categories methods for uncertainty representation in graph learning: direct inference, Bayesian representation learning, and Gaussian Processes.
IV-A Direct Inference
Direct inference is well-suited for semi-supervised node classification on graphs because a node’s neighboring nodes can be treated as new observed data, or evidence, allowing for direct updates to the node’s classification. One approach that directly applies Bayesian methods to infer the posterior probability distribution of nodes is as follows: First, a prior probability distribution is assigned to each node, which is then updated with evidence propagated from other nodes to obtain the posterior distribution. In [43], the node label is modeled as a categorical random variable as $P(\hat{y}_{i}=k|\boldsymbol{\theta})$, where $\hat{y}_{i}$ is the predicted label of a node and $\boldsymbol{\theta}$ is the parameter of the categorical distribution. Based on the assumption that connected nodes are likely to share a label, it is assumed that a neighbor of a node shares the same parameter $\boldsymbol{\theta}$ as the node itself. For a node $i$, this leads to the multinomial likelihood function for the labels of its neighbors: $P(\hat{N}_{i}|\boldsymbol{\theta})\propto\Pi_{k=1}^{K}\theta_{k}^{n_{ik}}$, where $\hat{N}_{i}$ represents the set of labeled neighbors of node $i$, and $n_{ik}$ is the number of neighbors of node $i$ with label $k$. The corresponding conjugate Dirichlet prior for the parameter $\boldsymbol{\theta}$ is given by: $P(\boldsymbol{\theta})\propto\Pi_{k=1}^{K}\theta_{k}^{\alpha_{k}-1}$, where $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\cdots,\alpha_{K})^{T}$ is the parameter of Dirichlet distribution. By combining these, the posterior distribution of $\boldsymbol{\theta}$ is: $P(\boldsymbol{\theta}|\hat{N}_{i})\propto\Pi_{k=1}^{K}\theta_{k}^{\alpha_{k}+n_{ik}-1}$.
Beyond this label-sharing assumption and label propagation, [44] introduced the propagation of multinomial messages to support both homophily and heterophily network effects. This was achieved by using compatibility matrices to represent the strength of connections between nodes. To handle high-dimensional features of nodes, [45] employed MLP encoding to compute node-level pseudo-counts, and then propagated them via PPR-based message passing. All three methods utilize Dirichlet priors and posteriors, with the main difference being the form of the evidence $n_{ik}$ that is propagated.
IV-B Bayesian Representation Learning
IV-B1 Bayesian Representation Learning Overview
Representation learning refers to the process of automatically learning meaningful and useful features or representations from raw data, typically through models like neural networks. The goal is to transform the data into a format that captures the underlying structure or patterns, making it more suitable for machine learning algorithms to process and extract insights. Unlike traditional deterministic methods, Bayesian representation learning models the uncertainty of both model parameters and learned representations, treating them as random variables with associated probability distributions instead of fixed values. In this survey, we focus on two prominent classes of Bayesian representation learning: variational autoencoders (VAEs) [46] and Bayesian neural networks (BNNs)  [47, 48].
VAEs. Autoencoders typically consist of two main components: an encoder and a decoder. The encoder, often denoted by a function such as $f_{\phi}(\mathbf{x})$, maps the input data $\mathbf{x}$ deterministically into a latent representation $\mathbf{z}$. The decoder, denoted by $g_{\theta}(\mathbf{z})$, then seeks to reconstruct the input data from this latent representation. The aim is to minimize the reconstruction error, typically measured by a loss function such as mean squared error for continuous input data or cross-entropy loss for binary input data. While traditional autoencoders output a fixed latent representation $\mathbf{z}$, VAEs extend this framework by introducing a variational encoder. This encoder learns the parameters of a probability distribution over the latent space, allowing $\mathbf{z}$ to be sampled from this distribution, thereby capturing the uncertainty in the representation.
A key technique in VAEs involves approximating the intractable posterior distribution $p_{\theta}(\mathbf{z}|\mathbf{x})$ using a simpler distribution $q_{\phi}(\mathbf{z}|\mathbf{x})$. To achieve this, the Evidence Lower Bound (ELBO) is maximized [46]:
$\mathcal{L}_{\text{ELBO}}(\theta,\phi;\mathbf{x})=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}[\log\,p_{\theta}(\mathbf{x}|\mathbf{z})]-\text{KL}[q_{\phi}(\mathbf{z}|\mathbf{x})\|p_{\theta}(\mathbf{z})].$ (14)
The first term is the expected log-likelihood and the second term is the KL divergence between the encoder’s distribution $q_{\phi}(\mathbf{z}|\mathbf{x})$ and $p_{\theta}(\mathbf{z})$, the prior distribution over the latent variables. VAEs capture uncertainty in the data by modeling the latent variables as distributions rather than fixed points. The variance of $q_{\phi}(\mathbf{z}|\mathbf{x})$ reflects this uncertainty. The KL divergence term in the ELBO balances fitting the data accurately while maintaining a prior over the latent space, embedding a principled approach to UQ in representation learning.
BNNs. Bayesian Neural Networks (BNNs) integrates the flexibility of neural networks with the probabilistic rigor of Bayesian inference. BNNs assign probability distributions to the neural network’s parameters $\boldsymbol{\theta}$, which encapsulates model uncertainty. This is a departure from traditional neural networks that only yield a single point estimate of parameters. During training, BNNs use Bayesian inference to update the prior distribution $P(\boldsymbol{\theta})$ with the likelihood $P(D|\boldsymbol{\theta})$, producing a posterior distribution that combines prior knowledge with observed data, as shown in Eq. (13). For predictions, BNNs leverage Bayesian model averaging, which involves integrating over the posterior distribution of the parameters to account for uncertainty. For a given test sample $\mathbf{x}^{*}$, the prediction thereby incorporates the model uncertainty, yielding the predictive distribution:
@@ IV-B1 Bayesian Representation Learning Overview
Due to the general intractability of the integral, various techniques for inferring $p(\boldsymbol{\theta}|D)$ have been proposed. [49] demonstrated that Monte Carlo dropout is equivalent to sampling from an approximate posterior of $\boldsymbol{\theta}$, allowing Eq. (15) to be simplified to the following Monte Carlo integral:
$p(y^{*}|\mathbf{x}^{*},D)\approx\frac{1}{S}\sum^{S}_{i=1}p(y^{*}|\mathbf{x}^{*},\boldsymbol{\theta}_{i}),$ (16)
where dropout is used to obtain $S$ weights $\boldsymbol{\theta}_{i}$. Given the challenges associated with exact inference, another widely-used technique for approximating the posterior distribution $p(\boldsymbol{\theta}|D)$ is variational inference (VI). The objective is to approximate a distribution $q_{\phi}(\boldsymbol{\theta})$ that is close to the posterior $p(\boldsymbol{\theta}|D)$, which is also achieved by maximizing the ELBO.
IV-B2 Bayesian Representation Learning for Graphs
@@
The need to represent the complexity of graph-structured data while quantifying uncertainty has driven the development of Bayesian graph representation learning. Unlike deterministic embeddings that provide fixed representations, Bayesian graph representation learning captures the stochastic nature of graphs, accounting for irregular structures, noisy features, and other uncertainties to produce more faithful representations. GNNs lay the groundwork by leveraging local neighborhood information, but their deterministic nature often falls short in capturing the inherent uncertainty of real-world graphs. To address these limitations, two prominent approaches have been developed: graph variational autoencoders (GVAEs) and Bayesian graph neural networks (BGNNs).
GVAEs. [50] proposed a novel framework for unsupervised learning on graph-structured data using variational autoencoders. The GCN based encoder transforms the input graph $G$ with nodes $V$ into a latent space $\mathbf{Z}$, where each node $i$ is mapped to a latent vector $\mathbf{z}_{i}$. The variational posterior $q(\mathbf{Z}|\mathbf{X},\mathbf{A})$ is approximated as a product of Gaussian distributions for each node. The model defines the posterior as $q(\mathbf{z}_{i}|\mathbf{x}_{i},\mathbf{A})=\mathcal{N}(\mathbf{z}_{i}|\boldsymbol{\mu}_{i},\text{diag}(\boldsymbol{\sigma}_{i}^{2}))$, where $\boldsymbol{\mu}$ and $\log\,\boldsymbol{\sigma}^{2}$ are the outputs of a GCN applied to the graph’s adjacency matrix $\mathbf{A}$ and node features $\mathbf{X}$. The objective function derived from the ELBO (Eq. 14) to be optimized is
$\mathcal{L}=\mathbb{E}_{q(\mathbf{Z}|\mathbf{X},\mathbf{A})}[\log\,p(\mathbf{A}|\mathbf{Z})]-\text{KL}[q(\mathbf{Z}|\mathbf{X},\mathbf{A})\|p(\mathbf{Z})],$ (17)
The term $\log\,p(\mathbf{A}|\mathbf{Z})$ represents the reconstruction loss, which encourages the model to accurately reconstruct the graph’s adjacency matrix. The KL divergence term $\text{KL}[q(\mathbf{Z}|\mathbf{X},\mathbf{A})\|p(\mathbf{Z})]$ acts as a regularizer, ensuring that the learned distribution does not deviate too far from the prior $p(\mathbf{Z})$.
The model in [50], while effective, has several limitations, including insufficient modeling of graph dependencies, inability to explain latent factors, and inadequate handling of isolated nodes, various methods have been introduced. To address these issues, various methods have been developed.
@@
[51] employed a hierarchical variational framework that enabled neighboring nodes to share information, thereby improving the modeling of the graph’s dependency structure. This approach allowed the propagation of graph structural information and distribution uncertainty, which are essential for capturing complex posterior distributions. To clearly explain what latent factors are and why they perform well, [52] introduced the Dirichlet Graph Variational Autoencoder (DGVAE) with graph cluster memberships as latent factors. This study connects VAE-based graph generation with balanced graph cut, providing a new perspective on how to enhance the internal mechanisms of VAEs for graph generation. [53] observed that existing autoencoders produce embeddings for isolated nodes that are close to zero, regardless of their content features. To address this issue, [53] proposed a novel Variational Graph Normalized AutoEncoder (VGNAE) that utilizes $L_{2}$-normalization to derive improved embeddings for isolated nodes. [54] proposed an ”iterative” GNN-based decoder to address the limitation of generating only node embeddings, as this is not conducive to sampling new graphs. In contrast to the node-level VGAE described earlier, [55] proposed a graph-level VAE [56], named GraphVAE, which modifies the encoder and decoder functions to work with graph-level latent representations, consequently facilitating the direct modeling of uncertainty within graphs.
@@
Capitalizing on the capabilities of graph variational autoencoders in representation and uncertainty modeling, several studies have incorporated them into GNNs for use in downstream tasks. [57] developed a novel hierarchical variational model that introduces additional latent random variables. This model jointly captures the hidden states of a Graph Recurrent Neural Network (GRNN) to effectively represent both topological changes and node attribute variations in dynamic graphs. To learn dynamic graph representations in hyperbolic space while also modeling uncertainty, [58] devised a hyperbolic graph variational autoencoder based on the proposed TGNN. This framework generates stochastic node representations from hyperbolic normal distributions.
@@
BGNNs. Another route lies in transferring general Bayesian neural networks to GNNs. These works consider model parameters, passed messages, etc., as distributions rather than deterministic values, and propose a series of BGNNs. To address the inability of standard GCNs to model uncertainty in graph, [59] treated the observed graph $G_{obs}$ as a sample from a random graph family. As a result, the posterior probability of node (or graph) labels can be computed by:
$p(\mathbf{Z}|\mathbf{Y}_{\mathcal{L}},\mathbf{X},G_{obs})=$ $\int p(\mathbf{Z}|\boldsymbol{\theta},G,\mathbf{X})p(\boldsymbol{\theta}|\mathbf{Y}_{\mathcal{L}},G,\mathbf{X})$
@@
Here $\boldsymbol{\theta}$ is a random variable representing the parameters of a Bayesian GCN over graph $G$, and $\lambda$ denotes the parameters that characterize a family of random graphs. $\mathbf{Y}_{\mathcal{L}}$ and $\mathbf{X}$ represent the training label set and the node feature matrix, respectively. $p(\mathbf{Z}|\boldsymbol{\theta},G,\mathbf{X})$ can be modeled using a categorical distribution by applying a softmax function to the output of the $L$-layer GCN $\mathbf{Z}=\mathbf{H}^{(L)}$. Due to the intractability of the integral in Eq. (IV-B2), the authors employ the following Monte Carlo approximation:
$p(\mathbf{Z}|\mathbf{Y}_{\mathcal{L}},\mathbf{X},G_{obs})\approx\frac{1}{N_{G}S}\sum_{i=1}^{N_{G}}\sum_{s=1}^{S}p(\mathbf{Z}|\boldsymbol{\theta}_{s,i},G_{i},\mathbf{X})\,.$ (19)
In this approximation, $N_{G}$ graphs $G_{i}$ are sampled from $p(G|\hat{\lambda})$, where $\hat{\lambda}$ is the parameter of the random graph model obtained through maximum a posteriori estimation $\hat{\lambda}=\arg\max_{\lambda}p(\lambda|G_{obs})$. $S$ weights $\boldsymbol{\theta}_{s_{i}}$ are drawn from $p(\boldsymbol{\theta}_{s,i}|\mathbf{Y}_{\mathcal{L}},G_{i},\mathbf{X})$ using Monte Carlo dropout. For the random graph model in Eq. (IV-B2), [59] utilized an assortative mixed membership stochastic block model (a-MMSBM)  [60, 61]. As an alternative, [62] proposed a node copying-based generative model, while [63] introduced a non-parametric posterior distribution to infer graph topology.
In addition to considering the observed graph as sampled from a random graph family, some studies also treat the inputs and other elements within the GNN as random variables, thereby adopting a Bayesian perspective to model uncertainty. Specifically, in these approaches, different components of the GNN are assigned probability distributions. Bayesian inference techniques are then employed to estimate the uncertainty associated with these components.
@@
To model the propagation of uncertainty in message-passing mechanisms, [64] treated messages as multivariate Gaussian variables and employed GNNs to predict their means. Recognizing that additional neighboring nodes provide more evidence and reduce uncertainty, [64] defined an uncertainty propagation mechanism in GNNs for predicting the covariance of Gaussian distributions, which is different from traditional message passing. The covariance ultimately provides the uncertainty of predictions. [65] viewed model parameters as random variables and then obtained both epistemic uncertainty and aleatoric uncertainty based on the uncertainty decomposition method in  [66, 67]. Additionally, this work incorporates uncertainty from evidence theory, namely vacuity due to lack of evidence and dissonance due to conflicting evidence, offering a richer representation of uncertainty. [68] considered the adjacency matrix $\mathbf{A}$ as a random variable. The prior distribution of the adjacency matrix $\mathbf{A}$ is given by:
$p(\mathbf{A})=\Pi_{ij}p(A_{ij}),\,\text{with}\,p(A_{ij})=\text{Bern}(A_{ij}|\rho_{ij}^{o}),$ (20)
@@
where $\text{Bern}(A_{ij}|\rho_{ij}^{o})$ is a Bernoulli distribution over $\rho_{ij}^{o}$. The variational posterior, which takes a form similar to the prior, is given by:
$q_{\phi}(\mathbf{A})=\Pi_{ij}q_{\phi}(A_{ij}),\,\text{with}\,q_{\phi}(A_{ij})=\text{Bern}(A_{ij}|\rho_{ij}),\,\rho_{ij}>0,$ (21)
where $\rho_{ij}$ are free parameters, and $\phi=\{\rho_{ij}\}$. By relaxing the discrete distribution to a continuous Concrete distribution [69, 70] and maximizing a ELBO, they are able to estimate the parameters $\phi$ of the posterior $q_{\phi}(\mathbf{A})$. [33] considered probabilistic links, noise in node features, and modeling errors, treating node features, links between nodes, and model parameters as random variables. The model propagates aleatoric uncertainty to the output by defining mechanisms for the means and variances of node embeddings in GNNs, while epistemic uncertainty is estimated using MC dropout.
IV-C Gaussian Process
IV-C1 Gaussian Process Overview

### v3

@@ IV Methods for Uncertainty Representation
TABLE II: Summary of Uncertainty Representation studies on graphs [^2] Reg., Cla., and Clu. are for Regression, Classification, and Clustering, respectively. NC and GC are for Node and Graph Classification, respectively. NRL and GRL are for Node and Graph Representation Learning, respectively. LP is for Link Prediction. STF is for Spatio-Temporal Forecasting. GG is for Graph Generation..
Methods Task Input Space Model Space
Node Feat. Graph Struc. GNN Msg. Model Params Latent Rep.
Bayesian Representation Learning
AEVB [56] DRL
BNNs [57] Reg. & Cla.
VGAE [58] NRL
GraphVAE [59] GRL
BGCNN [60] NC
BUP [61] NC
GKDE [62] NC
BGNN [44] NC
VGCN [63] NC
Gaussian Process
GGP [64] NC
UaGGP [65] NC
LNK [66] GRL
OOD Epistemic Uncertainty Representation
ADNCE [67] GC
WDRO [68] Clu.
DRGL [69] NC
PR-BCD [70] NC
Xu et al. [71] NC
DWNS [72] NC
RGIB [73] GC
GCC [74] NC & GC
Hu et al. [75] GC
Stochastic Differential Equations
LGNSDE [76] NC
GNSD [77] NC
DVGNN [78] STF
GraphGDP [79] GG
AGGDN [80] STF
TABLE III: Summary of Uncertainty Handling studies on graphs11footnotemark: 1.
Methods Task Post-hoc Non-parametric Input-Adaptive Topology-aware Note
Direct Inference
SocNL [81] NC Label Propagation
GPN [82] NC Evidence Propagation
NETCONF [83] NC Evidence Propagation
Out Of Distribution Techniques
DR-GNN [84] NC Distributionally Robust Optimization
Sadeghi, et al. [85] Reg. Distributionally Robust Optimization
WT-AWP [86] NC Adversarial Training
CAP [87] NC Adversarial Training
GraphDefense [88] NC Adversarial Training
BVAT [89] NC Adversarial Training
GraphAT [90] NC Adversarial Training
GraphCL [91] GC Self-Supervised Learning
SS-GCN [92] NC Self-Supervised Learning
GRACE [93] NC Self-Supervised Learning
Calibration
Histogram Binning [94] Cla. Binning-based
Isotonic Regression [95] Cla. Binning-based
Platt Scaling [96] Cla. Logit Modification
Label Smoothing [97] Cla. Label Modification
Focal Loss [98] Cla. Cross-entropy-based
CaGCN [99] NC Logit Modification
GATS [100] NC Logit Modification
RBS [101] NC Logit Modification
GCL [102] NC Cross-entropy-based
IN-N-OUT [37] LP Logit Modification
E-$\Delta$UQ [103] LP Anchoring-based
G-$\Delta$UQ [104] GC Anchoring-based
Conformal Prediction
Transductive CP [105] Reg. & Cla. Rely on Exchangeability
Inductive CP [106] Reg. & Cla. Rely on Exchangeability
Importance-weighted CP [107] Reg. & Cla. Beyond Exchangeability
CQR [108] Reg. Rely on Exchangeability
APS [109] Cla. Rely on Exchangeability
Robust Validation [110] Reg. & Cla. Beyond Exchangeability
Jurygcn [111] NC Beyond Exchangeability
Transductive DAPS [112] NC Rely on Exchangeability
Inductive NAPS [113] NC Beyond Exchangeability
{forest}
for tree= forked edges, grow’=0, draw=none, rounded corners, node options=align=center,, text width=2.7cm, anchor=center , [Uncertainty Quantification, fill=gray!10 ,rotate=90, parent [Uncertainty Representation Sec. IV, for tree=fill=teal!50, child [Bayesian Representation Learning Sec. IV-A, fill=teal!30, greatgrandchild [Graph Variational Autoencoders: VGAE [58], SIG-VAE [114], DGVAE [115], VGNAE [116], Graphite [117], GraphVAE [59], VGRNN [118], HVGAE [119], fill=teal!10, referenceblock] [Bayesian Graph Neural Networks: Bayesian GCNN [60], BGCN (node copying) [120], BGCN (non-parametric) [121], BUP [61], GKDE [62], VGCN [63], BGNN [44], fill=teal!10, referenceblock]] [Gaussian Process Sec. IV-B, fill=teal!30, greatgrandchild [GGP [64], UaGGP [65], LNK [66], GPGC [122], WGGP [123], GPG [124], DGPG [125], FT/WT-GP [126], GCLGP [127], GGPN [128], fill=teal!10, referenceblock]] [Outlier Representation Sec. IV-C, fill=teal!30, greatgrandchild [Dataset-level uncertainty estimation: ADNCE [67], WDRO [68], DRGL [69] , fill=teal!10, referenceblock] [Sample-level uncertainty representation: CAP [87], Xu et al. [71], DWNS [72], RGIB [73] , fill=teal!10, referenceblock] [Manually designed uncertainty sets: GraphCL [91], Hu et al. [75], GRACE [93] , fill=teal!10, referenceblock] ] [Stochastic Differential Equations Sec. IV-D, fill=teal!30, greatgrandchild [LGNSDE [76], GNSD [77], DVGNN [78], Graphgdp [79], AGGDN [80],fill=teal!10, referenceblock]] ] [Uncertainty Handling Sec. V, for tree=fill=orange!45,child [Bayesian Inference Techniques Sec. V-A, fill=orange!30, greatgrandchild [Exact: SocNL [81], NETCONF [83], GPN [82], fill=orange!10, referenceblock] [Approximate: VGAE [58], SIG-VAE [114], DGVAE [115], VGNAE [116], Graphite [117], GraphVAE [59], VGRNN [118], HVGAE [119],Bayesian GCNN [60], BGCN (node copying) [120], BGCN (non-parametric) [121], GKDE [62], VGCN [63], BGNN [44], GGP [64], UaGGP [65], LNK [66], WGGP [123], DGPG [125], FT/WT-GP [126], GCLGP [127], GGPN [128], LGNSDE [76], DVGNN [78], , fill=orange!10, referenceblock] ] [Out of Distribution Techniques Sec. V-B, fill=orange!30, greatgrandchild [Distributionally Robust Optimization: ADNCE [67], DR-GNN [84], WDRO [68], DRGL [69], Sadeghi, et al. [85] , fill=orange!10, referenceblock] [Adversarial Training: WT-AWP [86], CAP [87] PR-BCD [70], GraphDefense [88], DWNS [72], BVAT [89], GraphAT [90], RGIB [73] , fill=orange!10, referenceblock] [Self-Supervised Learning: GraphCL [91], GCC [74], SS-GCN [92], Hu et al. [75], GRACE [93] , fill=orange!10, referenceblock] ] [Calibration Sec. V-C, fill=orange!30, greatgrandchild [CaGCN [99], GCL [102], MCCNIFTY [129], IN-N-OUT [130], E-$\Delta$UQ [103], G-$\Delta$UQ [104],fill=orange!10, referenceblock]] [Conformal Prediction Sec. V-D, fill=orange!30, greatgrandchild [JuryGCN [111], DAPS [112], NAPS [113],fill=orange!10, referenceblock] ] ] ]
Figure 2: The taxonomy of uncertainty quantification methods in graph learning is organized into two main branches: Uncertainty Representation and Uncertainty Handling. Uncertainty Representation addresses how uncertainty is captured within data and models, including methods such as Bayesian Representation Learning, Gaussian Processes, and Outlier Representation. Uncertainty Handling focuses on methods for evaluating and mitigating uncertainty, encompassing Direct Inference, Out-of-Distribution Techniques, Calibration, and Conformal Prediction. Each leaf node highlights representative works and associated references for each method.
Understanding how uncertainty can be represented is fundamental for designing trustworthy graph learning algorithms. In this section, we delve into several distinct categories of methods for uncertainty representation, including bayesian representation learning, gaussian processes, stochastic differential equations (SDEs) and outlier representation.
IV-A Bayesian Representation Learning
IV-A1 Bayesian Representation Learning Overview
Bayesian methods in representation learning model uncertainty in both model parameters and learned representations. In this survey, we focus on two prominent classes: variational autoencoders (VAEs) [56] and Bayesian neural networks (BNNs)  [131, 132].
VAEs. VAEs introduce uncertainty representation by modeling the data‐generation process with fully probabilistic latent variables rather than a single deterministic code. Concretely, they assume a prior over latent variables $\mathbf{z}\sim p(\mathbf{z})$, commonly a standard normal, and define a likelihood $p_{\boldsymbol{\theta}}(\mathbf{x}\mid\mathbf{z})$, for example a Gaussian distribution whose learned variance captures the aleatoric uncertainty inherent in $\mathbf{x}$. To infer the latent representation for a given input $\mathbf{x}$, the encoder approximates the true posterior with a variational distribution $q_{\boldsymbol{\phi}}(\mathbf{z}\mid\mathbf{x})$, typically a Gaussian whose mean and variance depend on $\mathbf{x}$. Sampling $\mathbf{z}$ from $q_{\boldsymbol{\phi}}(\mathbf{z}\mid\mathbf{x})$ injects stochasticity into the encoding, quantifying uncertainty about which latent variable underlies the observation $\mathbf{x}$.
BNNs. Unlike deterministic neural networks that only provide point estimates, BNNs assign distributions to the parameters $\boldsymbol{\theta}$ to capture model uncertainty. During training, BNNs use Bayesian inference to update the prior distribution $p(\boldsymbol{\theta})$ with the likelihood $p(D|\boldsymbol{\theta})$, producing a posterior distribution that combines prior knowledge with observed data:
@@ IV-A1 Bayesian Representation Learning Overview
For predictions, BNNs leverage Bayesian model averaging, which involves integrating over the posterior distribution of the parameters to account for uncertainty. For a given test sample $\mathbf{x}^{\ast}$, the prediction incorporates the model uncertainty, yielding the predictive distribution:
@@
IV-A2 Bayesian Representation Learning for Graphs
@@
The need to represent the complexity of graph-structured data while quantifying uncertainty has driven the development of Bayesian graph representation learning. Unlike deterministic embeddings that provide fixed representations, Bayesian graph representation learning captures the stochastic nature of graphs, accounting for irregular structures, noisy features, and other uncertainties to produce more faithful representations. GNNs lay the groundwork by leveraging local neighborhood information, but their deterministic nature often falls short in capturing the inherent uncertainty of real-world graphs. In response, two prominent approaches have been developed: graph variational autoencoders (GVAEs) [58, 114, 115, 116, 117, 59, 119] and Bayesian graph neural networks (BGNNs) [60, 120, 121, 61, 62, 63, 44, 133].
GVAEs. VGAE [58] is one of the earliest works to apply VAEs to graph data. In VGAE, a GCN-based encoder processes the input graph $G=(V,E)$ with node features $\mathbf{X}$ and adjacency matrix $\mathbf{A}$, and maps each node $i\in V$ to a latent variable $\mathbf{z}_{i}$. Instead of producing a fixed embedding, the encoder defines a probabilistic representation by estimating a Gaussian distribution for each node:
$q_{\boldsymbol{\phi}}(\mathbf{z}_{i}\mid\mathbf{X},\mathbf{A})=\mathcal{N}\!\bigl{(}\mathbf{z}_{i}\mid\boldsymbol{\mu}_{i},\,\mathrm{diag}(\boldsymbol{\sigma}_{i}^{2})\bigr{)},$ (16)
where the mean vector $\boldsymbol{\mu}_{i}$ represents the central embedding and the variance $\boldsymbol{\sigma}_{i}^{2}$ quantifies the uncertainty associated with node $i$’s representation. On the decoder side, an inner product decoder is typically employed to reconstruct the graph structure from these uncertainty-aware latent representations.
@@
Although VGAE introduced probabilistic latent representations for graph, it still faces challenges in capturing complex dependency patterns, providing interpretable latent variables, and generating meaningful embeddings for isolated nodes. Consequently, various extensions have been proposed. Hasanzadeh et al. [114] propose a hierarchical variational framework that enables neighboring nodes to share information, thereby improving the modeling of the graph’s dependency structure. It allows the propagation of graph structural information and distribution uncertainty, which are essential for capturing complex posterior distributions. To clearly explain what latent factors are and why they perform well, Li et al. [115] introduces the Dirichlet Graph Variational Autoencoder (DGVAE) with graph cluster memberships as latent factors. They connect VAE-based graph generation with balanced graph cut, providing a new perspective on how to enhance the internal mechanisms of VAEs for graph generation. To avoid near-zero embeddings for isolated nodes generated by existing autoencoders, Variational Graph Normalized AutoEncoder (VGNAE) [116] utilizes $L_{2}$-normalization to derive improved embeddings for isolated nodes. Grover et al. [117] propose an “iterative” GNN-based decoder to progressively refine the reconstruction of graph adjacency matrices via low-rank approximations, enabling efficient and scalable unsupervised representation learning on large-scale graphs. Unlike the node-level methods described above, GraphVAE [59] modifies the encoder and decoder to work with latent representations of entire graphs. This approach focuses on modeling the uncertainty in graph-level representations, rather than in node-level ones.
@@
Capitalizing on the capabilities of GVAEs in representation and uncertainty modeling, several studies have extended them to model dynamic graphs more effectively. Hajiramezanali et al. [118] develop a novel hierarchical variational model that introduces additional latent random variables. This model jointly captures the hidden states of a Graph Recurrent Neural Network (GRNN) to effectively represent both topological changes and node attribute variations in dynamic graphs. To learn dynamic graph representations in hyperbolic space while also modeling uncertainty, TGNN [119] incorporating a hyperbolic graph variational autoencoder is proposed, which encodes stochastic nodes from hyperbolic normal distributions.
@@
BGNNs. A series of BGNN works consider model parameters, passed messages, and other key elements in graph learning as distributions rather than deterministic values. Bayesian GCNN [60] treats the observed graph $G_{obs}$ as a sample from an assortative mixed membership stochastic block model (a-MMSBM)  [134, 135]. As a result, the posterior probability of node (or graph) labels can be computed by:
$p(\mathbf{Y}^{*}|\mathbf{Y}^{*}_{\mathcal{L}},\mathbf{X},G_{obs})=$ $\int p(\mathbf{Z}|\boldsymbol{\theta},G,\mathbf{X})p(\boldsymbol{\theta}|\mathbf{Y}_{\mathcal{L}},G,\mathbf{X})$
@@
Here $\boldsymbol{\theta}$ is a random variable representing the parameters of a Bayesian GCN over graph $G$, and $\lambda$ denotes the parameters that characterize a family of random graphs. $\mathbf{Y}_{\mathcal{L}}$ and $\mathbf{X}$ represent the training label set and the node feature matrix, respectively. $p(\mathbf{Y}^{*}|\boldsymbol{\theta},G,\mathbf{X})$ can be modeled using a categorical distribution by applying a softmax function to the output of the $L$-layer GCN $\mathbf{H}^{(L)}$. As an alternative to a-MMSBM, BGCN (node copying) [120] uses a node copying-based generative model, while BGCN (non-parametric) [121] employs a non-parametric generative model to construct the posterior distribution of graph adjacency matrices.
In addition to treating the observed graph as a random variable, the following works also assign probability distributions to other GNN components, adopting a Bayesian approach to model uncertainty.
@@
To model the propagation of uncertainty in message-passing mechanisms, BUP [61] treats messages as multivariate Gaussian variables and employs GNNs to predict their means. Based on the assumption that additional neighboring nodes provide more evidence and reduce uncertainty, BUP defines an uncertainty propagation mechanism in GNNs for predicting the covariance of Gaussian distributions, which is different from traditional message passing. The covariance ultimately provides the uncertainty of predictions. GKDE [62] views model parameters as random variables and then obtains both EU and AU based on the uncertainty decomposition method in  [136, 137]. Additionally, GKDE incorporates uncertainty from evidence theory, namely vacuity resulting from lack of evidence and dissonance arising from conflicting evidence [138], offering a richer representation of uncertainty. [63] considered the adjacency matrix $\mathbf{A}$ as a random variable with prior distribution:
$p(\mathbf{A})=\prod_{ij}p(A_{ij}),p(A_{ij})=\text{Bern}(A_{ij}|\rho_{ij}^{o}),$ (18)
@@
where $\text{Bern}(A_{ij}|\rho_{ij}^{o})$ is a Bernoulli distribution over $\rho_{ij}^{o}$. The variational posterior is given by:
$q_{\phi}(\mathbf{A})=\prod_{ij}q_{\phi}(A_{ij}),q_{\phi}(A_{ij})=\text{Bern}(A_{ij}|\rho_{ij}),\,\rho_{ij}>0,$ (19)
where $\rho_{ij}$ are free parameters, and $\phi=\{\rho_{ij}\}_{i,j}$.
To separately quantify aleatoric and epistemic uncertainty, BGNN [44] treats probabilistic links, noise in node features and model parameters as random variables. Aleatoric uncertainty is propagated through mechanisms defining the means and variances of node embeddings, while epistemic uncertainty is estimated using MC dropout. In contrast to BGNN, Blei et al. [133] propose to decompose the total uncertainty based on the Bayesian inference of node labels. Total uncertainty is the posterior probability over the remaining unobserved node labels conditioned on acquiring the target node’s label; aleatoric uncertainty is the same posterior before label acquisition; and epistemic uncertainty is their ratio, reflecting the information gain from observing that label. In this decomposition, they find that leveraging epistemic uncertainty for node selection more effectively guides graph active learning.
IV-B Gaussian Process
IV-B1 Gaussian Process Overview